# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v1

@@ A.2.2 Contraction
Let $v_{1},\ldots,v_{n}$ be a basis of $V$ and write $v=\sum_{I}\lambda_{I}v_{I}$ where the sum is over multi-indices $I=(i_{1},\ldots,i_{r})$ and $v_{I}=v_{i_{1}}\wedge\ldots\wedge v_{i_{r}}$ are the basis elements of $\Lambda^{r}V$. Let $\eta_{1},\ldots,\eta_{n}$ be the dual basis defined by $\eta_{i}(v_{j})=1$ if $i=j$ and 00 otherwise. Suppose that $v$ is non-zero. Fix $I$ such that $\lambda_{I}\neq 0$ and let $i\in I$. Take $\omega=\eta_{i}$, we claim that $\omega\mathbin{\lrcorner}\,v\neq 0$. Indeed, if $I^{\prime}=I\setminus\{i\}$ then
@@
where the sum runs over multi-indices $J\neq I^{\prime}$ with $|J|=r-1$ which do not contain $i$ and $\tilde{\lambda}_{J}=\pm\lambda_{J\cup\{i\}}$. Conversely, if $\omega\neq 0$ then we can assume that $\omega=\eta_{1}$ and
@@ A.4 Distributions on $\mathbb{CP}^{n}$
A distribution on $\mathbb{CP}^{n}$ is a holomorphic vector subbundle $\mathcal{V}\subset T\mathbb{CP}^{n}$ defined outside a closed analytic subset $Z\subset\mathbb{CP}^{n}$ of complex codimension $\geq 2$. The rank $r$ of the subbundle is also called the rank of $\mathcal{V}$, we assume that $11\leq r\leq n-$ 1. The singular set of $\mathcal{V}$ is the smallest of all such $Z$’s, the regular set is the complement $U=\mathbb{CP}^{n}\setminus Z$.
@@ A.4.1 Homogeneous multivector fields on $\mathbb{C}^{n+1}$
[…]
@@
[…]
@@
[…]
@@
The case $\imath=r-d=0$ corresponds to genuine multivector fields on $\mathbb{CP}^{n}$ without any twisting. In this case the induced action $\mathbb{C}^{*}$ action on $V_{r+1,r+1}$ by scalar multiplication is trivial.
@@
Let $11\leq r\leq n-$ 1 and consider the linear subspace $M$ of $\mathbb{CP}^{n}$ of dimension $r-1$ given by the set of points $[x_{0}:\ldots:x_{n}]$ such that
@@
Let $p=n-r$ denote the codimension of a distribution. Then, an analogue of Proposition A.65, shows that there is a canonical identification between $\Omega^{p}_{\mathbb{P}^{n}}(n+1-\imath)$ (as in Remark A.60) and the vector space of $p$-forms on $\mathbb{C}^{n+1}$ with homogeneous coefficients of degree $d+1$ whose contraction with the Euler vector field is identically zero. In particular, codimension $1$ distributions on $\mathbb{CP}^{n}$ correspond to homogeneous $1$-forms
@@ A.4.2 Multivector fields tangent to hyperplanes
Let $X$ be a complex manifold and let $D\subset X$ be a smooth complex hypersurface. We say that an $r$-vector field $\mathbf{v}\in H^{0}(\Lambda^{r}TX)$ is tangent to $D$ at $x$ if $\mathbf{v}_{x}$ belongs to the linear subspace $\Lambda^{r}T_{x}D\subset\Lambda^{r}T_{x}X$ We say that $\mathbf{v}$ is tangent to $D$ if $\mathbf{v}$ is tangent to $D$ at all points $x\in D$.
@@
On the other hand, since at worst $I$ contains all the $n+1-k$ elements from $[n]\setminus I_{0}$
@@
By repeated applications of items (i) and (ii) of Lemma A.81, we can find $1$-forms $\omega_{i}\in W^{\circ}$ for $21\leq i\leq r-$ 2 such that
@@ Appendix B Essential and irreducible configurations
Let $P$ and $Q$ be linear subspaces of $\mathbb{CP}^{n}$ such that $P+Q=\mathbb{CP}^{n}$ and $P\cap Q\neq\emptyset$. Then $\dim P+\dim Q\geq n$
@@
$(\mathbf{e}_{\mathcal{B}})_{i}=\begin{cases}1&\text{ if }p_{i}\in\mathcal{B}\,,\\ 0&\text{ if }p_{i}\notin\mathcal{B}\,.\end{cases}$
@@
belongs to >$\mathbb{R}^{m}_{>0}$. It follows from Lemma B.11 (iii) and Lemma B.10 , that $\mathbf{a}_{0}\in C$. ∎
@@ Appendix C Miyaoka-Yau and Kähler-Einstein metrics
Let $(\mathcal{H},\mathbf{a})$ be a weighted arrangement of hyperplanes $H\subset\mathbb{CP}^{n}$ and let $(\mathbb{CP}^{n})^{\circ}$ be the arrangement complement. Let $\omega_{\operatorname{FS}}$ be the Fubini-Study metric on $\mathbb{CP}^{n}$ with $\omega_{\operatorname{FS}}\in 2\pi\cdot c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ and let ||$\cdot$|| ⋅ | be the usual Hermitian metric on $\mathcal{O}_{\mathbb{P}^{n}}(1)$ with curvature $-i\cdot\omega_{\operatorname{FS}}$. Consider the function
@@
where $\ell_{H}$ are linear functions on $z_{1},\ldots,z_{i}$ and ||$\cdot$|| ⋅ | is the usual absolute value on $\mathbb{C}$. By Fubini’s theorem it is enough to show that $f$ is locally in $L^{p}$ in a neighbourhood of the origin of the $\mathbb{C}^{i}$ factor transversal to $L$. At the same time, $U\setminus L\subset\Omega_{i-1}$ and by induction hypothesis $f\in L^{p}_{\operatorname{loc}}(U\setminus L)$. In particular, the integral of $|f|^{p}$ on a sphere about $0\in\mathbb{C}^{i}$ is finite. Taking spherical polar coordinates on $\mathbb{C}^{i}$ and by homogeneity of $f$, we see that $|f|^{p}$ is locally integrable around $0\in\mathbb{C}^{i}$ if and only if
[…]
@@
$N$ number of hyperplanes, $N=|\mathcal{H}|$

### v2

@@ A.2.2 Contraction
Let $v_{1},\ldots,v_{n}$ be a basis of $V$ and write $v=\sum_{I}\lambda_{I}v_{I}$ where the sum is over multi-indices $I=(i_{1},\ldots,i_{r})$ and $v_{I}=v_{i_{1}}\wedge\ldots\wedge v_{i_{r}}$ are the basis elements of $\Lambda^{r}V$. Let $\eta_{1},\ldots,\eta_{n}$ be the dual basis defined by $\eta_{i}(v_{j})=1$ if $i=j$ and 0 otherwise. Suppose that $v$ is non-zero. Fix $I$ such that $\lambda_{I}\neq 0$ and let $i\in I$. Take $\omega=\eta_{i}$, we claim that $\omega\mathbin{\lrcorner}\,v\neq 0$. Indeed, if $I^{\prime}=I\setminus\{i\}$ then
@@
where the sum runs over multi-indices $J\neq I^{\prime}$ with |J|=r1|J|=r-1 which do not contain $i$ and $\tilde{\lambda}_{J}=\pm\lambda_{J\cup\{i\}}$. Conversely, if $\omega\neq 0$ then we can assume that $\omega=\eta_{1}$ and
@@ A.4 Distributions on $\mathbb{CP}^{n}$
A distribution on $\mathbb{CP}^{n}$ is a holomorphic vector subbundle $\mathcal{V}\subset T\mathbb{CP}^{n}$ defined outside a closed analytic subset $Z\subset\mathbb{CP}^{n}$ of complex codimension $\geq 2$. The rank $r$ of the subbundle is also called the rank of $\mathcal{V}$, we assume that $1\leq r\leq n-1$. The singular set of $\mathcal{V}$ is the smallest of all such $Z$’s, the regular set is the complement $U=\mathbb{CP}^{n}\setminus Z$.
@@ A.4.1 Homogeneous multivector fields on $\mathbb{C}^{n+1}$
An exact sequence of vector bundles $0\to L\xrightarrow{\imath_{0}}V\xrightarrow{\pi}W\to 0$ with $\operatorname{rank}L=1$ determines an exact sequence
$0\to\Lambda^{r}W\otimes L\xrightarrow{\imath_{0}}\Lambda^{r+1}V\xrightarrow{\pi}\Lambda^{r+1}W\to 0,$
@@
$0\to\Lambda^{r}T\mathbb{CP}^{n}\xrightarrow{\imath_{0}}\Lambda^{r+1}E\xrightarrow{\pi}\Lambda^{r+1}T\mathbb{CP}^{n}\to 0\,$, (A.20)
@@
$0\to\Lambda^{r}T\mathbb{P}^{n}(-\imath)\xrightarrow{\imath_{0}}\Lambda^{r+1}E(-\imath)\xrightarrow{\pi}\Lambda^{r+1}T\mathbb{P}^{n}(-\imath)\to 0\,.$ (A.23)
Composing $\Lambda^{r+1}E(-\imath)\xrightarrow{\pi}\Lambda^{r+1}T\mathbb{CP}^{n}(-\imath)$ with the inclusion $\Lambda^{r+1}T\mathbb{CP}^{n}(-\imath)\xrightarrow{\imath_{0}}\Lambda^{r+2}E(-\imath)$ and using that $\imath_{0}\circ\pi=\wedge\,e$ we obtain an exact sequence
@@
The case $\imath=r-d=0$ corresponds to genuine multivector fields on $\mathbb{CP}^{n}$ - without any twisting. In this case the induced action $\mathbb{C}^{*}$ action on $V_{r+1,r+1}$ by scalar multiplication is trivial.
@@
Let $1\leq r\leq n-1$ and consider the linear subspace $M$ of $\mathbb{CP}^{n}$ of dimension r1r-1 given by the set of points $x_{0}:\ldots:x_{n}]$ such that
@@
Let p=nrp=n-r denote the codimension of a distribution. Then, an analogue of Proposition A.65, shows that there is a canonical identification between $\Omega^{p}_{\mathbb{P}^{n}}(n+1-\imath)$ (as in Remark A.60) and the vector space of $p$-forms on $\mathbb{C}^{n+1}$ with homogeneous coefficients of degree $d+1$ whose contraction with the Euler vector field is identically zero. In particular, codimension $1$ distributions on $\mathbb{CP}^{n}$ correspond to homogeneous $1$-forms
@@ A.4.2 Multivector fields tangent to hyperplanes
Let $X$ be a complex manifold and let $D\subset X$ be a smooth complex hypersurface. We say that an $r$-vector field $\mathbf{v}\in H^{0}(\Lambda^{r}TX)$ is tangent to $D$ at $x$ if $\mathbf{v}_{x}$ belongs to the linear subspace $\Lambda^{r}T_{x}D\subset\Lambda^{r}T_{x}X$ . We say that $\mathbf{v}$ is tangent to $D$ if $\mathbf{v}$ is tangent to $D$ at all points $x\in D$.
@@
On the other hand, since at worst II contains all the n+1kn+1-k elements from $n]\setminus I_{0}$
@@
By repeated applications of items (i) and (ii) of Lemma A.81, we can find $1$-forms $\omega_{i}\in W^{\circ}$ for $1\leq i\leq r-2$ such that
@@ Appendix B Essential and irreducible configurations
Let $P$ and $Q$ be linear subspaces of $\mathbb{CP}^{n}$ such that $P+Q=\mathbb{CP}^{n}$ and $P\cap Q\neq\emptyset$. Then $\dim P+\dim Q\geq n$ .
@@
[…]
@@
belongs to $\mathbb{R}^{m}_{>0}$. It follows from Lemma B.11 (iii) and Lemma B.10 , that $\mathbf{a}_{0}\in C$. ∎
@@ Appendix C Miyaoka-Yau and Kähler-Einstein metrics
Let $(\mathcal{H},\mathbf{a})$ be a weighted arrangement of hyperplanes $H\subset\mathbb{CP}^{n}$ and let $(\mathbb{CP}^{n})^{\circ}$ be the arrangement complement. Let $\omega_{\operatorname{FS}}$ be the Fubini-Study metric on $\mathbb{CP}^{n}$ with $\omega_{\operatorname{FS}}\in 2\pi\cdot c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ and let $|\cdot|$ be the usual Hermitian metric on $\mathcal{O}_{\mathbb{P}^{n}}(1)$ with curvature $i\cdot\omega_{\operatorname{FS}}$. Consider the function
@@
where $\ell_{H}$ are linear functions on $z_{1},\ldots,z_{i}$ and $|\cdot|$ is the usual absolute value on $\mathbb{C}$. By Fubini’s theorem it is enough to show that $f$ is locally in $L^{p}$ in a neighbourhood of the origin of the $\mathbb{C}^{i}$ factor transversal to $L$. At the same time, $U\setminus L\subset\Omega_{i-1}$ and by induction hypothesis $f\in L^{p}_{\operatorname{loc}}(U\setminus L)$. In particular, the integral of $|f|^{p}$ on a sphere about $0\in\mathbb{C}^{i}$ is finite. Taking spherical polar coordinates on $\mathbb{C}^{i}$ and by homogeneity of $f$, we see that $|f|^{p}$ is locally integrable around $0\in\mathbb{C}^{i}$ if and only if
$\int_{0}^{1}r^{-2\cdot p\cdot a_{L}}r^{2i-1}dr<\infty$ (C.3)
@@
NN number of hyperplanes, $N=|\mathcal{H}|$